OpenStudy (anonymous):

for the following system, if you isolate x in the second equation to use the substitution method, in the first equation? 3x +y =8 -x - y = -7

OpenStudy (anonymous):

-y+7 -y-7 y+7 y-7

OpenStudy (anonymous):

-y+7 because you would have to add y to the other side and then divide both sides by -1!

OpenStudy (anonymous):

dallas solved the system of equations below. what mistake did he make in his work? 3x + y=6 x - 2y =2 y=6-3x x-2 (6-3x)=2 x-12+6x=2 5x-12=2 +12 +12 5x-5 =14/5 x=2.8 3(2.8)+y=6 8.4+y=6 -8.4 -8.4 y=-2.4

OpenStudy (anonymous):

it should be 7x-12 on the 4th step!

OpenStudy (anonymous):

he combined like terms incorrectly. it should have been 7x instead od 5x.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

so x should equal 2 and y should equal 0 and yes you are correct!

OpenStudy (anonymous):

carrol high school has a total of 94 boys and girls who play sports. if the number of boys is 9 more than four times the number of girls, how many boys play sports at this high school?

OpenStudy (anonymous):

@Golferguy158

OpenStudy (anonymous):

your equation will be 94=4x +x +9 so you will get that there is 17 woman who play sports and 77 boys

OpenStudy (anonymous):

solve the following system. y= x +4 5x +y =16

OpenStudy (anonymous):

@Golferguy158